# Geometric foundations of thermodynamics in the quantum regime

### v1

@@ Abstract
In this work, we present a comprehensive geometrical formulation of quantum thermodynamics based on contact geometry and principal fiber bundles. The quantum thermodynamic state space is modeled as a contact manifold, with equilibrium Gibbs states forming a Legendrian submanifold that encodes the fundamental thermodynamic relations. A principal fiber bundle over the manifold of density operators distinguishes the quantum state structure from thermodynamic labels: its fibers represent non-equilibrium configurations, and their unique intersections with the equilibrium submanifold enforce thermodynamic consistency. Quasistatic processes correspond to minimizing geodesics under the Bures–Wasserstein metric, leading to minimal dissipation, while the divergence of geodesic length toward rank-deficient states geometrically derives the third law. Non-equilibrium extensions, formulated through pseudo-Riemannian metrics and connections on the principal bundle, introduce curvature-induced holonomy that quantifies irreversibility in cyclic processes. In this framework, the thermodynamic laws in the quantum regime emerge naturally as geometric consequences.
@@ I Introduction
Physics is geometry. Differential geometry [doCarmo1992Riemannian, ONeill1983SemiRiemannian, caja12_variedades, rudolph2012_1, rudolph2012_2] structures are fundamental to physical theories, from the symplectic geometry of classical mechanics [arnold1989mathematical] to the curved spacetime of general relativity [wald1984general] and the fiber bundle formulation of gauge theories [Bleecker1981Gauge, naber2011_1, naber2011_2, hamilton2017mathematical]. In thermodynamics [callen85_thermodynamics], contact geometry provides a natural and rigorous framework for equilibrium states and processes, as originally conceived by Gibbs and systematically developed in modern terms [bravetti15_contact, bravetti2017contact]. Here, the thermodynamic state space is a $(2n+1)$-dimensional contact manifold $(M,\eta)$, with the Gibbs 1-form $\eta$ encoding the first law, and equilibrium states forming a Legendrian submanifold $E$ where $\eta|_{E}=0$. Extending this geometric structure to the quantum regime [binder2018thermodynamics, strasberg2022quantum] presents significant conceptual and mathematical challenges. Recent works have explored quantum thermodynamics using “differential geometry” on the state space [scandi2019thermodynamic, anza2022geometric, tejero2025asymmetries, Bettmann2025Information]. However, there is not a proper geometrical formulation of the theory; just some processes analyzed with geometric perspective. A complete and unified geometric formulation of the full theory remains absent. Without a foundational geometric framework analogous to the classical case, analyses remain fragmented, and general principles are difficult to establish rigorously. Principal fiber bundles, central to gauge theories and topological phases in condensed matter [nakahara2018geometry, Bleecker1981Gauge], offer a powerful global structure for addressing these gaps. When applied to quantum thermodynamics, they furnish a principal bundle over the space of density operators, with fibers representing the thermodynamic coordinates associated with a fixed quantum state, thereby separating state evolution from thermodynamic variation and incorporating gauge-like symmetries.
@@
In this paper, we synthesize these elements into a comprehensive geometrical framework extended to the quantum regime. The quantum thermodynamic state space is indeed a contact manifold $\mathcal{M}$, where thermal equilibrium states form the Legendrian submanifold $\mathcal{E}\subset\mathcal{M}$, satisfying the quantum first law. We introduce a principal fiber bundle over the space of density operators, mapping thermodynamic coordinates to states. Fibers enable analysis of relaxation paths, with equilibrium as unique intersections with the equilibrium submanifold. The Bures-Wasserstein metric on such submanifold [bengtsson2006geometry, oostrum22_bures] defines geodesics as quasistatic processes minimizing dissipation. Boundary effects in finite dimensions lead to geodesic incompleteness, recovering the third law geometrically. Non-equilibrium phenomena are addressed naturally in the theory, via pseudo-Riemannian extensions of the metric, compatible with $\ker\eta$, and the structure of the fiber bundle. The principal connection decomposes the tangent bundle, with curvature inducing holonomy in cyclic processes, a geometric source of irreversibility akin to gauge holonomies. This geometrization not only generalizes classical geometric thermodynamics but also provides mathematical rigor for quantum thermodynamics.
@@ II Motivation: classical thermodynamics
Physics can be regarded as a collection of models that attempt to describe, always with some degree of idealization, specific aspects of the observable universe. Equilibrium thermodynamics exemplifies this as a robust physical theory centered on one of the most abstract idealizations: every system is assumed to be perpetually in thermodynamic equilibrium, implying that such a state would remain unchanged indefinitely. Paradoxically, this framework addresses phenomena like gas expansions, an out-of-equilibrium process. This is reconciled by positing that the evolution of the system proceeds exclusively through successive equilibrium states, necessitating an infinite duration for completion—a process termed quasistatic. Consequently, thermodynamic parameters vary smoothly, allowing well-defined derivatives with respect to an evolution parameter. These variations occur so gradually that they can be treated as infinitesimal, explaining why thermodynamic equations are typically expressed in differential forms. For instance, the first law is commonly written as $dU=TdS+\sum_{i}X_{i}dY_{i}+\sum_{j}\mu_{j}dN_{j}.$ The terms $dS$, $dY_{i}$, $dN_{j}$, for all $i,j$, can be interpreted as mere infinitesimals yielding finite integrals or, more insightfully, as differential 1-forms.
The latter perspective lends natural clarity to many thermodynamic concepts and results. Standard curricula on this topic usually distinguish state functions like entropy and internal energy—which depend solely on the system state when measured—from path-dependent quantities like heat or work. Thus, changes in internal energy or entropy are independent of the trajectory in thermodynamic parameter space, depending only on the endpoints, making them exact differentials. In contrast, integrals of heat or work generally vary with the path. A non-exact differential form may become exact when multiplied by an integrating factor, provided the form is closed. Assuming simply connected manifolds, closed and exact forms coincide. The prototypical example is heat: multiplying the infinitesimal amount $\delta Q$ by $1/T$ yields the exact differential $dS$.
Carathéodory’s theorem, often opaque in classical treatments, emerges straightforwardly geometrically: in any neighborhood of a thermodynamic state, certain states are inaccessible via quasistatic adiabatic processes. Geometrically, a process is a curve $\gamma:I\subseteq\mathbb{R}\to M$, where $M$ is the thermodynamic state space endowed with a differentiable manifold structure, as later explained. If $\delta Q\in\Lambda^{1}(M)$, the curve $\gamma$ is adiabatic when $\delta Q(\gamma^{\prime})=0$ everywhere. Since $dS=\delta Q/T$, it follows that $\delta Q=TdS$, so
@@
The geometrization of classical thermodynamics provides a powerful framework for analyzing thermodynamic systems using tools from differential geometry. This approach represents thermodynamic systems as manifolds endowed with specific geometric structures, enabling a rigorous description of equilibrium states and thermodynamic processes. Here, we introduce the thermodynamic state space and its contact structure, establishing the foundation for extending these concepts to quantum thermodynamics in subsequent sections. The thermodynamic state space is formalized as a $(2n+1)$-dimensional manifold $M$, where $n$ is the number of independent extensive variables characterizing the system. This manifold is equipped with a contact structure defined by a 1-form $\eta$, satisfying the non-integrability condition $\eta\wedge(d\eta)^{n}\neq 0$, where $\wedge$ denotes the wedge (also exterior or Grassmann) product and $(d\eta)^{n}$ is the $n$-fold exterior product of $d\eta$. The contact structure ensures that $\eta$ defines a maximally non-integrable distribution that captures the constraints of thermodynamic processes. In physical terms, the thermodynamic state space is endowed with the Gibbs 1-form, typically expressed for a simple thermodynamic system as
@@
Following Callen [callen85_thermodynamics] and Bravetti et al. [bravetti15_contact], we formulate classical thermodynamics through the following postulates, which provide a mathematically consistent foundation for the geometric approach.
@@
This geometric framework provides a robust foundation for extending thermodynamic concepts to quantum systems. In classical thermodynamics, the thermodynamic state space and its contact structure capture the interplay between extensive and intensive variables. In quantum thermodynamics, we aim to generalize this structure to account for quantum states, described by density operators on a Hilbert space, and their thermodynamic properties. The equilibrium submanifold is replaced by a quantum state space, and the Gibbs 1-form will be adapted to incorporate quantum mechanical constraints.
The third law of thermodynamics, which governs the asymptotic behavior of entropy as temperature approaches absolute zero, is not considered in this geometric analysis in the classical regime. However, in this quantum case, it will arise naturally from the mathematical structure of the theory.
@@ III Contact geometry
To establish the geometrical framework for quantum thermodynamics, we first review necessary concepts in contact geometry that underpin the thermodynamic state space and its extension to quantum systems. As stated, contact geometry provides a natural mathematical structure for describing thermodynamic equilibrium states and processes. This section introduces the essential definitions and results, setting the stage for their application to quantum thermodynamics in subsequent sections.
@@
Since $\eta\wedge(d\eta)^{n}\neq 0$, the differential $d\eta$ must have one-dimensional kernel, which is transverse to $\mathscr{D}$. This condition yields a unique vector field on the manifold, which governs the dynamics.
@@
In thermodynamics, Darboux theorem ensures that the thermodynamic state space can be locally described using coordinates that align the Gibbs 1-form with the contact structure. For example, for a simple system, coordinates can be chosen such that $\eta=dU-TdS+pdV$. This canonical form facilitates the analysis of thermodynamic relations and their extension to quantum systems.
@@ IV Connection to quantum thermodynamics
Let $\mathcal{H}$ be an $m$-dimensional complex Hilbert space, so that $\mathcal{H}\cong\mathbb{C}^{m}$ for $m<\infty$, and let $\mathscr{B}(\mathcal{H})$ be the set of bounded operators on $\mathcal{H}$. A quantum state is any $\rho\in\mathscr{B}(\mathcal{H})$ such that $\rho$ is self-adjoint, positive semidefinite and has unit trace.
See [gustafson03_mathematical, takhtadzhian08_quantum] for further details. In finite dimensions, $\mathscr{B}(\mathcal{H})\cong M_{m}(\mathbb{C})$. By virtue of this isomorphism, we identify states with unit trace self-adjoint matrices and introduce the following topological and differentiable structure.
@@
The positive definite states are those with full rank, i.e. $\mathrm{rank}(\rho)=m$, for all $\rho\in\mathcal{D}^{\circ}$.
@@
The manifold structure of $\mathcal{D}^{\circ}$ is trivial [caja12_variedades]. ∎
@@
The boundary $\partial\mathcal{D}$ contains all rank-deficient states, i.e. $\mathrm{rank}(\rho)<m$, including the pure states, i.e. states with rank $1$.
@@
Let $\mathcal{H}$ be a Hilbert space. An observable $A$ is a self-adjoint bounded operator on $\mathcal{H}$. For any quantum state $\rho\in\mathcal{D}$, the expectation value of $A$ in the state $\rho$ is defined as
@@
For later convenience, we introduce a finite set of observables $\boldsymbol{A}=\{A_{i}\in\mathscr{B}(\mathcal{H})\mid A_{i}\text{ observable},i=1,\ldots,n\}$. We associate to this set a collection of real-valued conjugate variables $\boldsymbol{\lambda}=\{\lambda_{i}\in\mathbb{R}\mid i=1,\ldots,n\}$, acting as smooth coordinates on the thermodynamic state space. Mimicking classical thermodynamics, the interest resides in defining a contact manifold for the quantum states, for them to be labeled by several coordinates by Theorem 1.
@@
where the smooth functions $\mu_{i}:\mathcal{M}\to\mathbb{R}$ satisfy the equilibrium consistency condition, i.e. given the embedding
@@
for all $\boldsymbol{\lambda}\in\mathbb{R}^{n}$, for $i=1,\ldots,n$.
@@
Differentiating the identity $S=\log Z+\sum_{i}\lambda_{i}a_{i}$ yields
@@
In thermodynamic equilibrium, i.e. on $\mathcal{E}$, the coordinates have physical interpretations. The coordinate $S$ is the von Neumann entropy $S=-\operatorname{\mathrm{tr}}(\rho\log\rho)$. Each $a_{i}=\operatorname{\mathrm{tr}}(A_{i}\rho)$ is the expectation value of the observable $A_{i}$. The conjugate parameters $\lambda_{i}$ play the role of intensive variables, e.g. inverse temperatures or chemical potentials. The contact relation $dS=\sum_{i}\lambda_{i}da_{i}$ is the quantum version of the classical first law in the entropy representation.
@@
Henceforth, thermodynamic entropy in the quantum regime is quantified by the von Neumann entropy, defined as $S(\rho)=-\operatorname{\mathrm{tr}}(\rho\ln\rho)$, for all $\rho\in\mathcal{D}^{\circ}$. Although the precise interpretation of quantum thermodynamic entropy remains a subject of active debate, for our purposes, the specific interpretation is immaterial, provided the entropy is a smooth, positive and concave function $S:\mathcal{D}^{\circ}\to\mathbb{R}$ on the manifold of density operators. These conditions ensure compatibility with the differential geometric framework developed herein.
The functions $\mu_{i}:\mathcal{M}\to\mathbb{R}$, $i=1,\dots,n$ allow $\Xi$ to be defined smoothly everywhere in $\mathcal{M}$, generalizing Gibbs states to non-equilibrium points. On the equilibrium submanifold $\mathcal{E}$, the consistency condition given by Eq. (18) enforces $\mu_{i}\left(S(\boldsymbol{\lambda}),\boldsymbol{a}(\boldsymbol{\lambda}),\boldsymbol{\lambda}\right)=\lambda_{i}$, for all $i=1,\ldots,n$ so that $\Xi\bigl(S(\boldsymbol{\lambda}),\boldsymbol{a}(\boldsymbol{\lambda}),\boldsymbol{\lambda}\bigr)=\rho_{\boldsymbol{\lambda}}$ recovers the standard Gibbs state as in Eq. (13). Off equilibrium, i.e. for points away from $\mathcal{E}$, the functions $\mu_{i}$ are not constrained by the Gibbs form and may be chosen flexibly to model non-equilibrium dynamics. For example, the correspondence $\mu_{i}=\lambda_{i}+f_{i}(S,\boldsymbol{a},\boldsymbol{\lambda})$, where each $f_{i}\in C^{\infty}(\mathcal{M})$ vanishes on $\mathcal{E}$, i.e. $f_{i}|_{\mathcal{E}}=0$, provides a valid extension, for all $i=1,\ldots,n$.
Remark 6.

### v2

@@ Abstract
In this work, we present a geometrical formulation of quantum thermodynamics based on contact geometry and principal fiber bundles. The quantum thermodynamic state space is modeled as a contact manifold, with equilibrium Gibbs states forming a Legendrian submanifold that encodes the fundamental thermodynamic relations. A principal fiber bundle over the manifold of density operators distinguishes the quantum state structure from thermodynamic labels: its fibers represent non-equilibrium configurations, and their unique intersections with the equilibrium submanifold enforce thermodynamic consistency. Quasistatic processes correspond to minimizing geodesics under the Bures–Wasserstein metric, leading to minimal dissipation, while the divergence of geodesic length toward rank-deficient states geometrically derives the unattainability aspect of the third law. Non-equilibrium extensions, formulated through pseudo-Riemannian metrics and connections on the principal bundle, introduce curvature-induced holonomy that quantifies a geometric source of irreversibility in cyclic processes. In this framework, the thermodynamic laws in the quantum regime emerge naturally as geometric consequences.
@@ I Introduction
Physics is geometry. Differential geometry [18, 35, 42, 38, 39] structures are fundamental to physical theories, from the symplectic geometry of classical mechanics [4] to the curved spacetime of general relativity [51] and the fiber bundle formulation of gauge theories [11, 30, 31, 22]. In thermodynamics [15], contact geometry provides a natural and rigorous framework for equilibrium states and processes, as originally conceived by Gibbs and systematically developed in modern terms [13, 14]. Here, the thermodynamic state space is a $(2n+1)$-dimensional contact manifold $(M,\eta)$, with the Gibbs 1-form $\eta$ encoding the first law, and equilibrium states forming a Legendrian submanifold $E$ where $\eta|_{E}=0$. Extending this geometric structure to the quantum regime [10, 44] presents significant conceptual and mathematical challenges.
Geometric methods have been increasingly applied to quantum thermodynamics, drawing from both differential geometry and information geometry to analyze dissipation, optimal processes, and fluctuations. In information geometry, the space of probability distributions (classical) or density operators (quantum) is endowed with a Riemannian metric, often the Fisher information metric or its quantum analogs like the quantum Fisher information [12, 2]. This framework has been used to link geometric quantities to thermodynamic concepts: for instance, the Fisher metric relates to thermalization speed and entropy production rates in stochastic thermodynamics [16, 25, 50, 29, 48, 7], and the quantum Fisher information bounds dissipation in finite-time quantum processes [17]. Other approaches emphasize Riemannian or pseudo-Riemannian metrics on the space of quantum states for optimization. For example, thermodynamic length—originally from classical thermodynamics [52, 40, 41]—has been generalized to quantum systems [43, 1, 7, 3].
Despite these advances, a complete and unified geometric formulation of the full theory is still lacking. In the absence of a geometric framework analogous to that of the classical case, existing analyses remain fragmented, and general principles are difficult to establish rigorously. The aforementioned works have investigated quantum thermodynamics using differential geometry, or tools drawn from it, such as representing processes as curves or defining thermodynamic lengths from metrics on the state space. However, these approaches do not constitute a proper geometric formulation of the theory; rather, they provide geometric perspectives on specific classes of processes. The geometric formulation proposed here builds on some of these ideas by reformulating the quantum state space within the language of contact geometry, extending classical contact thermodynamics, and incorporating the Bures–Wasserstein metric to describe quasistatic processes. These elements naturally connect with earlier uses of Riemannian metrics [40].
Principal fiber bundles, central to gauge theories and topological phases in condensed matter [33, 11], offer a powerful global structure for addressing these gaps. When applied to quantum thermodynamics, they furnish a principal bundle over the space of density operators, with fibers representing the thermodynamic coordinates associated with a fixed quantum state, thereby separating state evolution from thermodynamic variation and incorporating gauge-like symmetries.
@@
In this paper, we synthesize these elements into a geometrical framework extended to the quantum regime. The quantum thermodynamic state space is indeed a contact manifold $\mathcal{M}$, where thermal equilibrium states form the Legendrian submanifold $\mathcal{E}\subset\mathcal{M}$, satisfying the quantum first law. We introduce a principal fiber bundle over the space of density operators, mapping thermodynamic coordinates to states. Fibers enable analysis of relaxation paths, with equilibrium as unique intersections with the equilibrium submanifold. The Bures-Wasserstein metric on such submanifold [5, 36] defines geodesics as quasistatic processes minimizing dissipation. Boundary effects in finite dimensions lead to geodesic incompleteness, recovering the third law geometrically. Non-equilibrium phenomena are addressed naturally in the theory, via pseudo-Riemannian extensions of the metric, compatible with $\ker\eta$, and the structure of the fiber bundle. The principal connection decomposes the tangent bundle, with curvature inducing holonomy in cyclic processes, a geometric source of irreversibility akin to gauge holonomies.
With all these considerations, we construct a principal fiber bundle over the manifold of thermal states, separating quantum state structure from thermodynamic labels, enabling a unified treatment of equilibrium (Legendrian submanifold) and non-equilibrium (fiber paths). Then, the framework allows the understanding of the zeroth, first, second, and third laws as consequences of bundle properties, injectivity, contact non-integrability, and geodesic incompleteness, respectively. Finally, the curvature-induced holonomy acts as a gauge-like source of irreversibility in cyclic processes.
This geometrization not only generalizes classical geometric thermodynamics but also provides mathematical rigor for quantum thermodynamics, going beyond process-specific analyses by offering a geometric theory analogous to gauge theories. While the core framework developed here provides a geometric formulation for finite-dimensional quantum systems with a finite set of observables—where mathematical structures such as the principal bundle and Bures–Wasserstein metric are fully rigorous—we also discuss preliminary extensions to infinite-dimensional cases.
@@ II Motivation: classical thermodynamics
Physics can be regarded as a collection of models that attempt to describe, always with some degree of idealization, specific aspects of the observable universe. Equilibrium thermodynamics exemplifies this as a robust physical theory centered on one of the most abstract idealizations: every system is assumed to be perpetually in thermodynamic equilibrium, implying that such a state would remain unchanged indefinitely. Paradoxically, this framework addresses phenomena like gas expansions, an out-of-equilibrium process. This is reconciled by positing that the evolution of the system proceeds exclusively through successive equilibrium states, necessitating an infinite duration for completion—a process termed quasistatic. Consequently, thermodynamic parameters vary smoothly, allowing well-defined derivatives with respect to an evolution parameter. These variations occur so gradually that they can be treated as infinitesimal, explaining why thermodynamic equations are typically expressed in differential forms. For instance, the first law is commonly written as $dU=TdS+\sum_{i}X_{i}dY_{i}+\sum_{j}\mu_{j}dN_{j}.$ The terms $dS$, $dY_{i}$, $dN_{j}$, for all $i,j$, can be interpreted as mere infinitesimals leading to finite integrals or, more insightfully, as differential 1-forms.
The latter perspective lends natural clarity to many thermodynamic concepts and results. Standard curricula on this topic usually distinguish state functions like entropy and internal energy—which depend solely on the system state when measured—from path-dependent quantities like heat or work. Thus, changes in internal energy or entropy are independent of the trajectory in thermodynamic parameter space, depending only on the endpoints, making them exact differentials. In contrast, integrals of heat or work generally vary with the path. A non-exact differential form may become exact when multiplied by an integrating factor, provided the form is closed. Assuming simply connected manifolds, closed and exact forms coincide. The prototypical example is heat: by multiplying the infinitesimal amount $\delta Q$ by $1/T$ one obtains the exact differential $dS$.
Carathéodory’s theorem, often disregarded in classical treatments, emerges straightforwardly geometrically: in any neighborhood of a thermodynamic state, certain states are inaccessible via quasistatic adiabatic processes. Geometrically, a thermodynamic process is a curve $\gamma:I\subseteq\mathbb{R}\to M$, where $M$ is the thermodynamic state space endowed with a differentiable manifold structure, as later explained. If $\delta Q\in\Lambda^{1}(M)$, the curve $\gamma$ is adiabatic when $\delta Q(\gamma^{\prime})=0$ everywhere. Since $dS=\delta Q/T$, it follows that $\delta Q=TdS$, so
@@
The geometrization of classical thermodynamics provides a powerful framework for analyzing thermodynamic systems using tools from differential geometry. This approach represents thermodynamic systems as manifolds endowed with specific geometric structures, enabling a rigorous description of equilibrium states and thermodynamic processes. Here, we introduce the thermodynamic state space and its contact structure, establishing the foundation for extending these concepts to quantum thermodynamics in subsequent sections.
The thermodynamic state space is formalized as a $(2n+1)$-dimensional manifold $M$, where $n$ is the number of independent extensive variables characterizing the system. This manifold is equipped with a contact structure defined by a 1-form $\eta$, satisfying the non-integrability condition $\eta\wedge(d\eta)^{n}\neq 0$, where $\wedge$ denotes the wedge (also exterior or Grassmann) product and $(d\eta)^{n}$ is the $n$-fold exterior product of $d\eta$. The contact structure ensures that $\eta$ defines a maximally non-integrable distribution that captures the constraints of thermodynamic processes. In physical terms, the thermodynamic state space is endowed with the Gibbs 1-form, typically expressed for a simple thermodynamic system as
@@
Following Callen [15] and Bravetti et al. [13], we formulate classical thermodynamics through the following postulates, which provide a mathematically consistent foundation for the geometric approach.
@@
This geometric perspective provides a robust foundation for extending thermodynamic concepts to quantum systems. In classical thermodynamics, the thermodynamic state space and its contact structure capture the interplay between extensive and intensive variables. In the quantum case, we aim to generalize this structure to account for quantum states, described by density operators on a Hilbert space, and their thermodynamic properties. The equilibrium submanifold is replaced by a quantum state space, and the Gibbs 1-form will be adapted to incorporate quantum mechanical constraints.
The third law of thermodynamics, which governs the asymptotic behavior of entropy as temperature approaches absolute zero, is not considered in this geometric analysis in the classical regime. However, in this quantum case, a version of the third law will arise naturally from the mathematical structure of the theory.
@@ III Contact geometry
To establish the geometrical framework for quantum thermodynamics, we first review the necessary concepts in contact geometry that underpin the thermodynamic state space and its extension to quantum systems. As stated, contact geometry provides a natural mathematical structure for describing thermodynamic equilibrium states and processes. This section introduces the essential definitions and results, setting the stage for their application to quantum thermodynamics in subsequent sections.
@@
Since $\eta\wedge(d\eta)^{n}\neq 0$, the differential $d\eta$ must have one-dimensional kernel, which is transverse to $\mathscr{D}$. This condition leads to the existence of a unique vector field on the manifold, which governs the dynamics.
@@
In thermodynamics, Darboux’s theorem ensures that the thermodynamic state space can be locally described using coordinates that align the Gibbs 1-form with the contact structure. For example, for a simple system, coordinates can be chosen such that $\eta=dU-TdS+pdV$. This canonical form facilitates the analysis of thermodynamic relations and their extension to quantum systems.
The elements of contact geometry collectively provide a rigorous mathematical encoding of the intrinsic constraints and processes in thermodynamics. Physically speaking, the contact form $\eta$ embodies the non-integrable nature of the first law. The contact distribution $\ker(\eta)$ corresponds to directions in the state space where reversible transformations can occur. The Reeb vector field $\mathcal{R}_{\eta}$, which corresponds to $\partial_{S}$ in Darboux’s coordinates, defines directions transverse to the equilibrium submanifolds.
@@ IV Connection to quantum thermodynamics
Let $\mathcal{H}$ be an $m$-dimensional complex Hilbert space, so that $\mathcal{H}\cong\mathbb{C}^{m}$ for $m<\infty$, and let $\mathscr{B}(\mathcal{H})$ be the set of bounded operators on $\mathcal{H}$. A quantum state is any $\rho\in\mathscr{B}(\mathcal{H})$ such that $\rho$ is self-adjoint, positive semidefinite, and has unit trace.
See [21, 46] for further details. In finite dimensions, $\mathscr{B}(\mathcal{H})\cong M_{m}(\mathbb{C})$. By virtue of this isomorphism, we identify states with unit trace self-adjoint matrices and introduce the following topological and differentiable structure.
@@
The positive definite states are those with full rank, i.e., $\mathrm{rank}(\rho)=m$, for all $\rho\in\mathcal{D}^{\circ}$.
@@
The manifold structure of $\mathcal{D}^{\circ}$ is trivial [42]. ∎
@@
The boundary $\partial\mathcal{D}$ contains all rank-deficient states, i.e., $\mathrm{rank}(\rho)<m$, including the pure states, i.e., states with rank $1$.
@@
Let $\mathcal{H}$ be a Hilbert space. An observable $A$ is a self-adjoint operator on $\mathcal{H}$. For any quantum state $\rho\in\mathcal{D}$, the expectation value of $A$ in the state $\rho$ is defined as
@@
For later convenience, we introduce a finite set of observables $\boldsymbol{A}=\{A_{i}\in\mathscr{B}(\mathcal{H})\mid A_{i}\text{ observable},i=1,\ldots,n\}$. We associate with this set a collection of real-valued conjugate variables $\boldsymbol{\lambda}=\{\lambda_{i}\in\mathbb{R}\mid i=1,\ldots,n\}$, acting as smooth coordinates on the thermodynamic state space. Mimicking classical thermodynamics, the interest resides in defining a contact manifold for the quantum states, for them to be labeled by several coordinates by Theorem 1.
@@
where the smooth functions $\mu_{i}:\mathcal{M}\to\mathbb{R}$ satisfy the equilibrium consistency condition, i.e., given the embedding
@@
for all $\boldsymbol{\lambda}\in\mathbb{R}^{n}$, for $i=1,\ldots,n$. Additionally, for fixed values of $S$ and $\boldsymbol{a}$, and also $c_{1},\dots,c_{n}\in\mathbb{R}$, the system
$\mu_{1}(S,\boldsymbol{a},\boldsymbol{\lambda})=c_{1},\hskip 36.0pt\ldots\hskip 36.0pt\mu_{n}(S,\boldsymbol{a},\boldsymbol{\lambda})=c_{n},$ (19)
has a unique solution $\lambda_{1},\dots,\lambda_{n}\in\mathbb{R}$. This condition will be called joint injectivity of the $\mu_{i}$ functions.
@@
Differentiating the identity $S=\log Z+\sum_{i}\lambda_{i}a_{i}$
@@
In thermodynamic equilibrium, i.e., on $\mathcal{E}$, the coordinates have physical interpretations. The coordinate $S$ is the von Neumann entropy $S=-\operatorname{\mathrm{tr}}(\rho\log\rho)$. Each $a_{i}=\operatorname{\mathrm{tr}}(A_{i}\rho)$ is the expectation value of the observable $A_{i}$. The conjugate parameters $\lambda_{i}$ play the role of intensive variables, e.g., inverse temperatures or chemical potentials. The contact relation $dS=\sum_{i}\lambda_{i}da_{i}$ is the quantum version of the classical first law in the entropy representation.
@@
Henceforth, thermodynamic entropy in the quantum regime is quantified by the von Neumann entropy, defined as $S(\rho)=-\operatorname{\mathrm{tr}}(\rho\log\rho)$, for all $\rho\in\mathcal{D}^{\circ}$. Although the precise interpretation of quantum thermodynamic entropy remains a subject of active debate, for our purposes, the specific interpretation is immaterial, provided the entropy is a smooth, positive, and concave function $S:\mathcal{D}^{\circ}\to\mathbb{R}$ on the manifold of density operators. These conditions ensure compatibility with the differential geometric framework developed herein.
The functions $\mu_{i}:\mathcal{M}\to\mathbb{R}$, $i=1,\dots,n$ allow $\Xi$ to be defined smoothly everywhere in $\mathcal{M}$, generalizing Gibbs states to non-equilibrium points. On the equilibrium submanifold $\mathcal{E}$, the consistency condition given by Eq. (18) enforces $\mu_{i}\left(S(\boldsymbol{\lambda}),\boldsymbol{a}(\boldsymbol{\lambda}),\boldsymbol{\lambda}\right)=\lambda_{i}$, for all $i=1,\ldots,n$ so that $\Xi\bigl(S(\boldsymbol{\lambda}),\boldsymbol{a}(\boldsymbol{\lambda}),\boldsymbol{\lambda}\bigr)=\rho_{\boldsymbol{\lambda}}$ recovers the standard Gibbs state as in Eq. (13). Off equilibrium, i.e., for points away from $\mathcal{E}$, the functions $\mu_{i}$ are not constrained by the Gibbs form and may be chosen flexibly to model non-equilibrium dynamics. For example, the correspondence $\mu_{i}=\lambda_{i}+f_{i}(S,\boldsymbol{a},\boldsymbol{\lambda})$, where each $f_{i}\in C^{\infty}(\mathcal{M})$ vanishes on $\mathcal{E}$, i.e., $f_{i}|_{\mathcal{E}}=0$, provides a valid extension, for all $i=1,\ldots,n$.